OpenStudy (anonymous):

what does this graph look like? y = 2(5x)

OpenStudy (anonymous):

@andu1854

OpenStudy (unklerhaukus):

is that \[y=2\times(5x)\]or\[y=2^{(5x)}\]?

OpenStudy (anonymous):

y=2(5^x)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

simply plug in the numbers. so you would plug in x to find y. ex. u put in x=1. then y=(5*1). Do you understand it?

OpenStudy (anonymous):

or if it is y=2(5^x) it would be 2(5^1). Or are you in algebra 2 and trying to find out the shape from the problem?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Can you just show me the graph? lol

OpenStudy (anonymous):

@blahblah_who_cares

OpenStudy (anonymous):



OpenStudy (anonymous):

If you came to this website for people to do your hw then there should be no reason to help you.